OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

A) 12x + 4y = -12 15x = -30 B) 8x + 4y = -4 14x = -10 C) 6x - 8y = -36 -6y = -42 D) 6x - y = -15 3y = 9

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

First, find the solution to your system. Then plug it into the other options.

OpenStudy (anonymous):

so solve these? 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

I think I did it, I got A. @SolomonZelman